OpenStudy (xxstudyingstuffxx):

How do you write ​0.000013​ in scientific notation?

OpenStudy (xxstudyingstuffxx):

Fan and medal

OpenStudy (narissa):

Havent done this is a while but i think its...1.3 x 10-5.

OpenStudy (rcp031):

scientific notation? means is a special way of writing numbers: How to Do it. When the number is smaller than 1, the decimal point has to move to the right, so the power of 10 is negative. Example: 0.0055 is written 5.5 × 10^-3 Because 0.0055 = 5.5 × 0.001 = 5.5 × 10^-3

OpenStudy (narissa):

1.3 x 10^-5

OpenStudy (xxstudyingstuffxx):

So the answer is \[1.3 \times 10^{-5}\]

OpenStudy (xxstudyingstuffxx):

?

OpenStudy (narissa):

I believe so :) Maybe get a second opinion I'm not very good at math.

OpenStudy (xxstudyingstuffxx):

Thanks. I would say I am bad at math, but I got moved up a grade because I was too good at math for 7th grade.

OpenStudy (rcp031):

u can also right 13*10^-6 or it is equivalent to saying (1/10^6) of 13

OpenStudy (narissa):

Ah thats so awesome I wish I got moved up a grade but my grades go up and down and back up to much lol

OpenStudy (xxstudyingstuffxx):

Thanks guys!